OpenStudy (anonymous):

help. can you explan how to do this and do it as an example for me? Solve the literal equation of the stated variable. A=piR^2h, for r

OpenStudy (anonymous):

the first step is to find R^2

OpenStudy (anonymous):

so far i have \[\frac{ A }{ h}=piR^2\]

OpenStudy (anonymous):

you still have pi "stuck" on R^2 ,to get rid of it divide both sides by pi

OpenStudy (anonymous):

A=piR^2h ___ ___ pi pi

OpenStudy (anonymous):